# Libra: Leveraging Temporal Images for Biomedical Radiology Analysis

### v1

@@ Abstract
Radiology report generation (RRG) is a challenging task, as it requires a thorough understanding of medical images, integration of multiple temporal inputs, and accurate report generation. Effective interpretation of medical images, such as chest X-rays (CXRs), demands sophisticated visual-language reasoning to map visual findings to structured reports. Recent studies have shown that multimodal large language models (MLLMs) can acquire multimodal capabilities by aligning with pre-trained vision encoders. However, current approaches predominantly focus on single-image analysis or utilise rule-based symbolic processing to handle multiple images, thereby overlooking the essential temporal information derived from comparing current images with prior ones. To overcome this critical limitation, we introduce Libra, a temporal-aware MLLM tailored for CXR report generation using temporal images. Libra integrates a radiology-specific image encoder with a MLLM and utilises a novel Temporal Alignment Connector to capture and synthesise temporal information of images across different time points with unprecedented precision. Extensive experiments show that Libra achieves new state-of-the-art performance among the same parameter scale MLLMs for RRG tasks on the MIMIC-CXR. Specifically, Libra improves the RadCliQ metric by 12.9% and makes substantial gains across all lexical metrics compared to previous models111Our code is open-sourced at https://github.com/X-iZhang/Libra..
@@ 1 Introduction
Radiology reports are a key component of biomedical radiology analysis, designed to enhance clarity and efficiency in medical communication. These structured reports are typically divided into distinct sections, such as Findings, Impression, Indication, Technique, Comparison, and History (Ganeshan et al., 2018). Serving as the primary medium for radiologists to convey their findings and conclusions from imaging studies like chest X-rays (CXRs), radiology reports play a crucial role in guiding diagnostic and therapeutic decisions for various diseases (Najjar, 2023). However, generating these reports manually is a complex and time-consuming task. Automating radiology report generation (RRG) presents a valuable opportunity to boost radiologist productivity, improve communication, and alleviate burnout (Zhang et al., 2020b). Despite its potential, RRG is a challenging task due to the intricate nature of medical imaging and the need for precise and accurate documentation.
Although the advancements in paradigm models like LLaVA (Liu et al., 2023) and InstructBLIP (Dai et al., 2023), they still exhibit a significant performance gap when applied to RRG tasks. In particular, their effectiveness diminishes in biomedical contexts due to the substantial differences between biomedical and general image-text pairs (Tu et al., 2023; Saab et al., 2024). Consequently, these models often perform at a surface level, akin to a layperson-level understanding, when dealing with specialised medical imaging tasks. Recent advancements have explored continued pre-training of general-purpose foundation models for medical tasks, but these models still fall short in addressing the complexities of medical image analysis (Li et al., 2023a; Chaves et al., 2024; Park et al., 2024).
Harnessing the full potential of multimodal large language models (MLLMs) for understanding and reasoning over multimodal data in the biomedical domain poses a significant challenge, primarily due to the granularity and specificity required in medical image analysis (Wang et al., 2023a). Preliminary attempts to apply MLLMs to RRG have largely concentrated on generating reports from a single image, often neglecting the crucial temporal relationships between images taken at different time points (Zhang et al., 2024b). The MIMIC-CXR Database (Johnson et al., 2019a) reveals that 67% of patients underwent at least two studies at different time intervals, underscoring the necessity of incorporating temporal context in clinical practice. This highlights a critical gap in existing models, which are not equipped to handle the dynamic progression of medical conditions reflected in sequential imaging studies.
In standard clinical practice, radiologists rely heavily on comparing current imaging results with previous studies to assess temporal changes. However, many models inadvertently introduce spurious references to prior examinations that do not exist during inference, leading to inaccurate reports. Recent advancements222Detailed related work is discussed in Appx. A.1, and our research objectives are explained in Appx. A.2., such as MedVersa (Zhou et al., 2024) and MAIRA-2 (Bannur et al., 2024), can process multiple images simultaneously. These models insert visual tokens from different images at specific points within the textual input, hinging on the LLM to interpret reference information. Additionally, LLaVA-Rad (Chaves et al., 2024) utilised GPT-4V (OpenAI et al., 2024) for radiology report extraction and augmentation, eliminating hallucinations associated with prior image references in the dataset. Despite these efforts, existing models still lack integrated mechanisms for perceiving temporal information within their architectures, which restricts the model’s ability to achieve the temporal awareness required for comprehensive reporting. Meanwhile, existing MLLMs typically utilise visual embeddings from the last or penultimate layer of the image encoder (Chen et al., 2023a; Zhang et al., 2024a), which can only capture the global characteristics of the image. However, RRG tasks require not only identifying findings but also capturing subtle details such as severity, extent, and the progression of findings (Sloan et al., 2024). Such fine-grained details are difficult to represent using embeddings from any single layer of the image encoder (Jiang et al., 2024). Therefore, for the high-granularity demands of RRG tasks, relying solely on representation from a single layer limits the quality of report generation. To tackle these limitations, we aim to enhance the temporal awareness of MLLMs for RRG task by tackling two main challenges:
It is non-trivial to design effective structures in MLLMs to handle prior study citations across various time points in the RRG task.
@@
The scarcity of effective connectors in MLLMs capable of handling the high-granularity requirements of downstream tasks.
@@
To overcome these challenges, we propose Libra (Leveraging Temporal Images for Biomedical Radiology Analysis), a novel framework designed to incorporate temporal change information into the RRG task. Libra employs a pre-trained visual transformer, RAD-DINO (Pérez-García et al., 2024), as the image encoder to generate robust image features, which are then refined using a new adapter specifically designed for the temporal awareness, before being fed into the medical large language model (LLM), Meditron (Chen et al., 2023b). Through a two-stage training strategy, Libra demonstrates the potential of MLLMs for specialised radiology applications. Our modular approach integrates state-of-the-art open-source pre-trained models for image and text while using a temporal-aware adapter to align these modalities with the text embedding space. Extensive experiments show that Libra substantially outperforms existing MLLMs in RRG tasks, demonstrating superior cross-modal understanding and reasoning capabilities. Specifically, our contributions include:
We present Libra, the temporal-aware MLLM capable of capturing temporal changes and overcoming the challenge of handling prior study citations, setting a new state-of-the-art performance in RRG tasks on the MIMIC-CXR dataset among MLLMs of the same parameter scale.
@@
We designed the Temporal Alignment Connector (TAC) with two components: the Layerwise Feature Extractor (LFE), which extracts high-granularity image feature embedding from the encoder, and the Temporal Fusion Module (TFM), which integrates temporal references from prior studies. This enhances Libra’s ability to capture and utilise temporal information in the RRG task.
2 Libra: Being Aware of Temporal Changes in Chest X-rays Report Generation
@@ 2.1 Model architecture
Libra model follows the common architecture of MLLMs, such as LLaVA, which includes an image encoder, a text decoder (i.e., an LLM pre-trained based on text-only data) and an adapter module for mapping visual representations into text space. In this work, we utilise a frozen biomedical image encoder, RAD-DINO (Pérez-García et al., 2024), a visual transformer extensively pre-trained on medical scans using the DINOv2 image-only self-supervised learning approach (Oquab et al., 2024). The LLM is deployed by Meditron-7B (Chen et al., 2023b), which builds on Llama-2 (Touvron et al., 2023) and further pre-trained on a specialised medical corpus. However, for the adapter module, we specifically designed a Temporal Alignment Connector (TAC) to capture and integrate temporal information across two images from different time points, as shown in Figure 1. The TAC bridges the image encoder and LLM. It consists of two components: the Layerwise Feature Extractor (LFE), which extracts high-granularity image representations, and the Temporal Fusion Module (TFM), which integrates temporal references from prior studies. This design enables Libra to effectively manage temporal data and enhances its ability to generate accurate and coherent radiology reports.
2.2 Temporal Alignment Connector (TAC): New Adapter Design
To address the challenges of integrating temporal information and aligning high-granularity image features in RRG tasks, we developed the TAC module. It bridges the gap between visual features extracted from multiple temporal snapshots of a patient’s CXRs and the language model. The TAC consists of two sub-modules: the LFE, which is to extract the layerwise representation of images, and the TFM, which is to capture the temporal-aware representation between two images.
2.2.1 Layerwise Feature Extractor (LFE)
@@
To leverage abundant feature information in a pre-trained image encoder, we extract all image patch token features from each hidden layer for a given input image. Specifically, RAD-DINO has 12 hidden layers and processes 518 × 518 images into 14 × 14 patches, generating 1,369 patch token sequences per hidden layer for each image. We do not use the token. Following this, we obtain patch embeddings of the same dimension from each layer, denoted as , where is the number of patch tokens and is the embedding dimension of the image encoder. Then, we combine them as , where is the number of hidden layers. Drawing from VGG (Simonyan & Zisserman, 2015), we employ a progressive compression strategy to reduce the layer dimension, ensuring that the LFE captures the most relevant features. Initially, we utilise Squeeze-and-Excitation (SE) Networks (Hu et al., 2019), which construct informative features by integrating both spatial and channel-wise information within local receptive fields at each layer. The SE block is applied to obtain calibrated feature representations, using GELU (Hendrycks & Gimpel, 2023) as the activation function. Next, we employ a specialised pointwise convolution module to align the feature spaces across different layers, using a depthwise 2D convolution with filters and stride of , without bias. The dimensions are gradually reduced to better align with the text in the LLM. The compressed features are represented as , where is the original layer number and is the layer number after compression. Following the size-reduction pattern of convolutional layers in VGG, the image representations are compressed according to 333Since RAD-DINO has 12 hidden layers, the prime factorisation chain provides the factors as .. Through three stages of progressive compression, we obtain the final representation of image patch features. We define the function to project image features into the same dimension:
where , . The LFE independently processes each image to produce a unified representation for temporal alignment, as (a) in Figure 1.
2.2.2 Temporal Fusion Module (TFM)
The fusion module is inspired by the transformer decoder. It utilises prior images as auxiliary information to weight the current image, generating a refined, temporal-aware representation suitable for the LLM’s latent space. The TFM uses an attention mechanism to dynamically learn relationships between image pairs, enabling the model to adapt to various temporal changes.
Prior Image Prefix Bias: The dataset contains samples with and without a prior image. The patch token features of the current and previous images processed by the LFE are denoted as and . If no true prior image exists, we set , effectively using the current image as a dummy prior. To differentiate this case, we introduce a trainable bias . Following the attention mechanism’s scaling techniques for adjusting hidden space degrees of freedom with a chi-square distribution (Vaswani, 2017; Wang et al., 2018a), we apply an exponent of to the cosine similarity, where is the hidden dimension of the LLM, to obtain a weighted bias addition to :
This nonlinear scaling amplifies higher similarity values, effectively modulating the influence of prior image features. When no true prior image is available, the high similarity score ensures that the effect of the dummy prior is adequately represented. This adjustment prevents samples with a dummy prior image from undergoing redundant rounds of parallel multi-head self-attention during subsequent propagation through the transformer blocks, as shown in Figure 1.
Transformer Block: Our transformer block includes a multi-head cross-attention sub-layer, multi-head self-attention sub-layers, and two multi-layer perceptron (MLP) sub-layers. As shown in Figure 1 (b). The paired are processed with layer normalization and residual connections:
where is a simple neural network composed of two fully connected layers with GELU as the activation function. After that, the features are processed through , a straightforward neural network consisting of four fully connected layers with the same activation function, but with hidden dimensions matching those of the LLM. We define our fusion module function as .
@@ 2.2.2 Temporal Fusion Module (TFM)
Subsequently, becomes the input sequence that LLM can comprehend, where is the number of patch tokens and is the hidden dimension of LLM. The final output is a refined feature representation that encapsulates the temporal evolution of the patient’s condition, making it suitable for the language model to generate accurate and contextually aware radiology reports.
@@ 2.3 Prompt Design
To enhance Libra’s ability to perceive temporal changes and integrate medical information in the RRG task, we designed a comprehensive prompting strategy comprising a system prompt and detailed clinical prompts, as shown in Figure 1. The system prompt is crafted to enable the LLM to recognise temporal variations: “The assistant specialised in comparing Chest X-ray images, identifying differences, and noting temporal changes.” Following this, additional sections of the report, such as {Indication}, {History}, {Comparison}, and {Technique}, are incorporated into the clinical instructions (e.g., in Appx. B.4). The clinical prompt to Libra is: “Provide a detailed description of the findings in the radiology image. Following clinical context: .” If no clinical instructions are available, the prompt defaults to: “Provide a detailed description of the findings in the radiology image.” After tokenizing and embedding the prompt and the answer, we insert the image patch tokens at the specified location, typically between the system prompt and the clinical prompts.
@@ 2.4 Temporal-aware Training
In this study, we focus on frontal-view images, either posterior-anterior (PA) or anterior-posterior (AP), and the Findings sections of radiology reports, as they contain the most direct clinical observations. We adopt a two-stage training strategy inspired by recent advancements in MLLM fine-tuning techniques (McKinzie et al., 2024). This proposed approach aims to progressively teach the model essential skills, including visual feature alignment and temporal information extraction. Libra’s training process involves two stages: temporal feature alignment and downstream task fine-tuning.
In the first stage, the visual encoder and LLM weights are frozen, and the TAC is trained with Findings and Impression generation, along with CXR-related VQA tasks to extract high-quality image representations and capture temporal changes. In the second stage, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2021) to fine-tune the pre-trained LLM on the Findings section generation task, keeping the visual encoder and connector weights frozen. Unlike traditional full fine-tuning, LoRA achieves comparable performance with significantly lower training costs. The detailed training configuration, including learning rate schedules and model parameters, is provided in Appx. B.1.

### v2

@@ Abstract
Radiology report generation (RRG) requires advanced medical image analysis, effective temporal reasoning, and accurate text generation. While multimodal large language models (MLLMs) align with pre-trained vision encoders to enhance visual-language understanding, most existing methods rely on single-image analysis or rule-based heuristics to process multiple images, failing to fully leverage temporal information in multi-modal medical datasets. In this paper, we introduce Libra, a temporal-aware MLLM tailored for chest X-ray report generation. Libra combines a radiology-specific image encoder with a novel Temporal Alignment Connector (TAC), designed to accurately capture and integrate temporal differences between paired current and prior images. Extensive experiments on the MIMIC-CXR dataset demonstrate that Libra establishes a new state-of-the-art benchmark among similarly scaled MLLMs, setting new standards in both clinical relevance and lexical accuracy.
Libra: Leveraging Temporal Images for Biomedical Radiology Analysis
Xi Zhang1,2, Zaiqiao Meng1,2, Jake Lever1,2 & Edmond S. L. Ho
1Information Retrieval Group, 2AI4BioMed Lab
School of Computing Science
University of Glasgow
{X.Zhang.6}@research.gla.ac.uk
{Zaiqiao.Meng,Jake.Lever,Shu-Lim.Ho}@glasgow.ac.uk
https://x-izhang.github.io/Libra
@@ 1 Introduction
Radiology reports are critical for biomedical radiology analysis, offering structured summaries of imaging studies such as chest X-rays (CXRs). Commonly divided into sections like Findings, Impression, Indication, Technique, Comparison, and History (Ganeshan et al., 2018), these reports guide diagnostic and therapeutic decisions (Najjar, 2023). However, manually generating such reports is both complex and time-consuming. Automating radiology report generation (RRG) holds great promise for alleviating radiologist burnout, increasing efficiency, and improving communication (Zhang et al., 2020b). Despite this, the intricate nature of medical imaging demands precise and detailed documentation, making RRG a challenging task.
Recent advances in Multimodal Large Language Models (MLLMs), such as LLaVA (Liu et al., 2023) and InstructBLIP (Dai et al., 2023), have demonstrated potential in vision-language tasks. However, their performance diminishes in specialised biomedical contexts due to the significant domain shift between general-purpose and medical image-text data (Tu et al., 2023). These models often lack fine-grained detail for medical imaging tasks, resulting in surface-level understanding akin to layperson interpretations. While continued pre-training on medical datasets and domain-specific fine-tuning (e.g., LLaVA-Med (Li et al., 2023a), MedBLIP (Chen et al., 2023b)) improve performance, they still cannot fully capture the complexities of medical image analysis (Xiao et al., 2024).
One critical gap in the current MLLM-based approaches is their limited ability to incorporate temporal context, which is pivotal in clinical practice. Radiologists routinely compare current imaging results with prior studies to identify temporal changes, a process crucial for understanding disease progression and guiding treatment decisions. Indeed, the MIMIC-CXR database (Johnson et al., 2019b) reveals that 67% of patients underwent at least two studies at different time intervals, underscoring the necessity of temporal reasoning. However, most MLLMs designed for RRG tasks focus on single-image analysis, neglecting this temporal dimension (Zhang et al., 2024c). As illustrated in Figure 1, MAIRA-1 (Hyland et al., 2024) introduces hallucinated prior references in single-image analysis and misinterprets temporal changes when integrating prior studies.
Although recent models111Detailed related work is discussed in Appx. A, and our research objectives are explained in Appx. B. like MedVersa (Zhou et al., 2024) and MAIRA-2 (Bannur et al., 2024) have introduced multi-image processing, they do not explicitly model or extract temporal differences. Instead, they rely on inserting visual tokens from different studies at specific points within textual inputs and delegate the reasoning task to the LLM. Similarly, Banerjee et al. (2024) and Chaves et al. (2024) leverage GPT-4V (OpenAI et al., 2024) to eliminate hallucinated references to prior studies in the dataset but lacks dedicated mechanisms for modelling temporal progression. Additionally, existing MLLMs often rely on embeddings from the last or penultimate layer of the image encoder (Chen et al., 2023a; Zhang et al., 2024a), primarily capturing global features. However, RRG tasks require fine-grained details222E.g., severity and temporal progression of findings. (Sloan et al., 2024), which a single-layer embedding often cannot fully represent (Jiang et al., 2024). To tackle these limitations, we enhance MLLM temporal awareness for RRG tasks by addressing two main challenges:
Designing robust MLLM architectures that seamlessly handle prior study references in RRG.
@@
The scarcity of effective feature alignment projectors in MLLMs capable of handling the high-granularity requirements of downstream tasks.
@@
To overcome these gaps, we propose Libra (Leveraging Temporal Images for Biomedical Radiology Analysis), a novel temporal-aware framework tailored for RRG tasks. Libra employs a pre-trained visual transformer encoder, RAD-DINO (Pérez-García et al., 2024), to generate robust image features, which are then refined using a new projector crafted for the temporal awareness, before being fed into the medical large language model (LLM), Meditron (Chen et al., 2023c). Through a two-stage training strategy, Libra aligns temporal visual features with the text embedding space, improving temporal coherence in RRG.
Our modular approach integrates state-of-the-art open-source pre-trained models for medical image and text processing while introducing a dedicated temporal-aware adapter to align visual and textual modalities within the embedding space. This paper makes the following contributions:
Libra, a temporal-aware MLLM designed to model temporal references and mitigate temporal hallucinations in RRG tasks.
@@
Temporal Alignment Connector (TAC), comprising the Layerwise Feature Extractor (LFE) and Temporal Fusion Module (TFM), which extracts high-granularity image features from multiple encoder layers and integrates temporal references from the prior study when available.
Extensive evaluation on the MIMIC-CXR dataset, achieving state-of-the-art results on average among similarly scaled MLLMs, with case analysis illustrating Libra’s architectural benefits.
2 Libra
@@ 2.1 Model Architecture
Our Libra model follows the standard architecture of MLLMs, such as LLaVA (Liu et al., 2023), comprising an image encoder, a text decoder and a connector module to map visual features into the text embedding space. Figure 2 shows the overall architecture of Libra. Specifically, we utilise a frozen biomedical image encoder, i.e. RAD-DINO (Pérez-García et al., 2024), a visual transformer extensively pre-trained on medical scans using the DINOv2 image-only self-supervised learning approach (Oquab et al., 2024). The text encoder is deployed by Meditron-7B (Chen et al., 2023c), which builds on Llama-2 and is further pre-trained on specialised medical corpora.
To effectively connect the image encoder and LLM, we design a novel Temporal Alignment Connector (TAC) tailored to capture and integrate temporal information from paired images taken at different time points. Meanwhile, when no prior image is available, we employ a dummy prior image, which is simply a copy of the current image, to mitigate spurious references to nonexistent scans, as shown in Figure 2 (bottom). This design enables Libra to effectively manage temporal data (e.g., stable, improved, worsening) and enhances its ability to generate accurate and coherent radiology reports.
2.2 Temporal Alignment Connector
To address the challenges of integrating temporal information and aligning high-granularity visual features for RRG tasks, TAC bridges the image encoder and the LLM. It processes visual features from two temporal snapshots to produce a unified representation sensitive to temporal changes. As shown in Figure 2 (right), TAC includes two key components: the Layerwise Feature Extractor, which extracts high-granularity image representations, and the Temporal Fusion Module, which integrates temporal references from the prior study.
2.2.1 Layerwise Feature Extractor
To leverage abundant image feature representations encoded by a pre-trained image encoder, we extract image patch token features of all the hidden layers for a given pair of input images. By default, the RAD-DINO image encoder (Pérez-García et al., 2024) has 12 hidden layers and processes 518 × 518 input images into 14 × 14 patches, generating 1,369 patch token sequences per hidden layer. Rather than relying on a single global feature token (e.g., the token), we collect same-dimensional patch embeddings from each layer per image, denoted as , where is the number of patch tokens and is the embedding dimension of the image encoder.
Then, these embeddings are concatenated across all layers as , where is the number of hidden layers. Drawing from the progressive compression strategy in VGG (Simonyan and Zisserman, 2015), our Layerwise Feature Extractor (LFE) reduces dimensionality across layers while preserving critical information. First, we utilise Squeeze-and-Excitation (SE) Networks (Hu et al., 2019), which construct informative features by integrating both spatial and channel-wise information within local receptive fields at each layer. The SE block is applied to obtain calibrated feature representations, using GELU (Hendrycks and Gimpel, 2023) as the activation function.
@@ 2.2.1 Layerwise Feature Extractor
Next, we employ a specialised pointwise convolution module to align the feature spaces across different layers, using a depthwise 2D convolution with filters and stride of 1, without bias. The compressed features are represented as , where is the original layer number and is the layer number after compression. Following the size-reduction pattern of convolutional layers in VGG, the image features are compressed according to 333Since RAD-DINO has hidden layers, the prime factorisation chain provides the factors as .. Through three stages of progressive compression, we obtain the final patch-level representation:
For simplicity, we use to denote the above three stages of compression, which project a given input image into its feature representation of the fixed dimension, :
By progressively refining each image’s representations through multiple stages, the LFE generates a unified and compact feature set suitable for temporal alignment. This design ensures that both high-granularity and global context are retained, as illustrated in (a) of Figure 2.
2.2.2 Temporal Fusion Module
The Temporal Fusion Module (TFM) is inspired by the transformer decoder and is designed to integrate temporal information by leveraging prior images as auxiliary context. It takes as input a paired set of compressed features from both the current and prior images, denoted as and , respectively, which are obtained after processing through the LFE. The temporal fusion process is defined as:
@@ 2.2.2 Temporal Fusion Module
where TFM learns to weigh the current image using prior image features, refining the representation to enhance temporal awareness. The resulting feature sequence, , serves as the input to the LLM, where is the number of patch tokens and is the hidden dimension of the LLM. This process encapsulates the temporal evolution of the patient’s condition, allowing the language model to generate accurate and contextually aware radiology reports.
Prior Image Prefix Bias
The dataset contains samples with and without a prior image. When a prior image is not available, we set . However, this “dummy prior image” is indistinguishable from a true prior in raw features. To differentiate it, we add a trainable bias, as .
Following the attention scaling techniques for adjusting hidden space degrees of freedom with a chi-square distribution (Vaswani, 2017), a nonlinear scaling function amplifies higher similarity values. The cosine similarity between the current and prior images is scaled with an exponent of , where is the hidden dimension of the LLM:
This nonlinear scaling emphasises high similarity values, modulating the influence of prior image features. When no prior image is available, the high similarity score ensures that the effect of the dummy prior is adequately represented. This adjustment prevents samples with a dummy prior image from undergoing redundant rounds of parallel multi-head self-attention during subsequent propagation through the transformer blocks, in Figure 2.
Transformer Block
The Transformer Block in TFM follows the standard Transformer design but is optimized for handling temporal image pairs. It consists of multi-head self-attention (), multi-head cross-attention (), and two multi-layer perceptron () sub-layers. As illustrated in (b) of Figure 2. The paired are processed with layer normalization () and residual connections:
where is a simple neural network composed of two fully connected layers with GELU as the activation function. After that, the features are processed through , a straightforward neural network consisting of four fully connected layers with the same activation function, but with hidden dimensions matching those of the LLM.
@@ 2.3 Prompt Design
To enhance Libra’s ability to perceive temporal changes and integrate medical information in RRG, we design a structured prompting strategy, consisting of a system prompt and a clinical prompt, as shown in Figure 2 (left). The system prompt enables the LLM to recognise temporal variations, while standard report sections (Indication History, Comparison, Technique) are integrated into the clinical prompt (see Appx. C for a detailed example).
The full prompt is: “Provide a detailed description of the findings in the radiology image. Following clinical context:.” There are datasets, e.g. MIMIC-CXR (Johnson et al., 2019b), where the report sections are unavailable. For these datasets, we set the prompt as follows: “Provide a detailed description of the findings in the radiology image.” After tokenising and embedding prompts, the refined image features () are inserted between the system prompt and clinical prompts.
@@ 2.4 Temporal-aware Training
Libra focuses on frontal-view images, either posterior-anterior (PA) or anterior-posterior (AP), and targets the Findings sections of RRG, as these contain the most direct clinical observations. It employs a two-stage training strategy, inspired by recent MLLM fine-tuning techniques (McKinzie et al., 2024), to progressively learn visual feature alignment and temporal information extraction.
Temporal Feature Alignment
In the first stage, the visual encoder and LLM weights are frozen, while the TAC is trained. This stage focuses on Findings and Impression generation from paired images and performing CXR-related visual question answering (VQA) tasks to extract high-quality image representations and capture temporal changes.
Downstream Task Fine-tuning
In the second stage, we apply Low-Rank Adaptation (LoRA) (Hu et al., 2021) to fine-tune the LLM on the Findings generation task, while keeping the visual encoder and TAC weights frozen. LoRA achieves performance comparable to full fine-tuning at a substantially lower computational cost. The detailed training configuration, including learning rate schedules and hyperparameters, is provided in Appx. D.
| Lexical: | |||||||
| ROUGE-L | 27.6 | 21.5 | 13.2 | 27.5 | 30.6 | 28.9 | 36.2 (+18.3%) |
| BLEU-1 | 35.4 | 16.9 | 16.4 | 32.3 | 38.1 | 39.2 | 51.2 (+30.6%) |
| BLEU-4 | 14.9 | 4.7 | 17.8 | 11.5 | 15.4 | 14.2 | 24.3 (+36.5%) |
| METEOR | 35.3 | – | – | – | – | 33.3 | 48.7 (+38.0%) |
| Clinical: | |||||||
| RadGraph-F1 | 19.1 | – | – | 26.7 | – | 24.3 | 32.4 (+21.3%) |
| RG | 23.8 | 20.5 | 13.2 | – | 29.4 | 29.6 | 36.9 (+25.0%) |
| RadCliQ() | 3.30 | – | – | – | – | 3.10 | 2.76 (+11.0%) |
| CheXbert vector | 36.9 | – | – | – | – | 44.0 | 46.3 (+5.2%) |
| CheXpert-F1: | |||||||
| Micro-F1-14 | 42.7 | 39.3 | 35.5 | 53.6 | 57.3 | 55.7 | 55.3 (-3.4%) |
| Macro-F1-14 | 26.9 | 24.7 | 20.4 | 39.8 | 39.5 | 38.6 | 40.2 (+1.1%) |
| Micro-F1-5 | 43.9 | 41.2 | 25.8 | 57.9 | 57.4 | 56.0 | 58.9 (+1.8%) |
| Macro-F1-5 | 36.3 | 34.5 | 19.6 | 51.6 | 47.7 | 47.7 | 52.6 (+2.0%) |